Isotopes of oxygen with mass number less than 16 undergo beta plus emission.  Assuming an equi molar mixture of O14 and O15, find a) the ratio of nucleide at the end of 1 Hr and b) at what time will the ratio of O14  nuclei to O15 nuclei be equal to 0.25.?  Given that t1/2 for O14=71sec and t1/2 for O15= 124 sec.

Grade:11

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Kiran,

O14 ----> N14         λ1

O15-----> N15         λ2

λ1=0.693/71=0.00976

λ2=0.693/124=0.00589

[A]01=[A]02

λ1* t=ln{ [A]01/[A]t1}----1

λ2* t=ln{ [A]02/[A]t2}-----2

After 1 hr=60*60 sec

a)from 1 and 2 :  eλ1*t/eλ2*t=[A]t2/[A]t1=e(λ1-λ2)*t=e0.00387*60*60=e13.932

b)[A]t2/[A]t1=e(λ1-λ2)*t=e0.00387*t=1/0.25

   0.00387*t = ln 4=2.303*2*0.3010 =1.386

   t= 1.386/0.00387 =358.14 sec